250.0 mL of a 0.500 M HCl solution needs to be made from concentrated HCl. What volume of the concentrated solution is needed if its molarity is 12.0 M?



1
Expert's answer
2021-03-26T05:41:31-0400

Proportion 1:

0.500 mol – 1000 mL

x – 250 mL

"x = \frac{0.500 \times 250}{1000} = 0.125 \;mol"

Proportion 2:

12.0 mol – 1000 mL

0.125 mol – y mL

"y = \frac{0.125 \times 1000}{12.0} = 10.41 \;mL"

Answer: 10.41 mL